OpenStudy (anonymous):

Solve: (x)/(x-2) + (1)/(x-4) = (2)/(x^2 - 6x + 8)

OpenStudy (lgbasallote):

multiply with LCM this time it's (x-2)(x-4) that is equal to x^2 -6x +8 x(x-4) + x-2 = 2 x^2 - 4x + x -2 -2 =0 x^2 - 3x -4 =0 (x-4)(x+1) = 0 x = 4 x=-1

OpenStudy (anonymous):

What is the LCD? Let's start there.

OpenStudy (anonymous):

Least Common Denominator?

OpenStudy (anonymous):

That's right - first we want to identify the Least Common Denominator.

OpenStudy (anonymous):

Can you tell what it is?

OpenStudy (anonymous):

Well I don't really know how to tell D:

OpenStudy (anonymous):

Alright - I'll explain how to find it. :)

OpenStudy (anonymous):

When you are asked to solve rational equations like this, the first thing you want to do is factor everything.

OpenStudy (anonymous):

Factor them by what? LCD?

OpenStudy (anonymous):

Factor each polynomial, all the numerators and denominators. I'll show you:

OpenStudy (anonymous):

This is what you were given: \[\frac{x}{x-2} + \frac{1}{x-4} = \frac{2}{x^2 - 6x + 8} \]Now, first identify 3 fractions. Each fraction has a numerator and denominator, but only one of these pieces can be factored. The denominator on the right-most fraction.

OpenStudy (anonymous):

To factor it, we're looking specifically at this: \[x^2-6x+8\]Here, you want to ask yourself, what two numbers multiply to +8 and add to -6.

OpenStudy (anonymous):

is it because it is a trinomial?

OpenStudy (anonymous):

That's right! :) Thanks for all the medals, by the way.

OpenStudy (anonymous):

Can you see how it would factor?

OpenStudy (anonymous):

It would factor like this: We're looking for numbers that multiply to 8 and add to -6. So the numbers we are looking for are -4 and -2. So this trinomial would factor like this: (x-4)(x-2) And you can always FOIL those back out to check to see that you get what you started with.

OpenStudy (anonymous):

I was eating late dinner sorry lol and your welcome. thanks for the tutor! (:

OpenStudy (anonymous):

Happy to help! You ready for the next step?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Alright. :)

OpenStudy (anonymous):

you make it easier than my online class really is :P

OpenStudy (anonymous):

So we can rewrite the original question in factored form, like this:\[\frac{x}{x-2} + \frac{1}{x-4} = \frac{2}{(x-4)(x-2)} \]

OpenStudy (anonymous):

The reason we do this is because we are eventually going to multiply both sides of this by the LCD in order to completely clear the fractions. But, we have to know what factors all the pieces are made from in order to do that first.

OpenStudy (anonymous):

So at this point we can see that the LCD is: (x-4)(x-2)

OpenStudy (anonymous):

alright so what do i do with the LCD

OpenStudy (anonymous):

From here, you're going to multiply each fraction by the LCD over 1 - like this:

OpenStudy (anonymous):

\[\frac{(x-4)(x-2)}{1}\frac{x}{x-2} + \frac{(x-4)(x-2)}{1}\frac{1}{x-4} = \frac{(x-4)(x-2)}{1}\frac{2}{(x-4)(x-2)} \]

OpenStudy (anonymous):

It's a long process, but now comes the good part - you get to cancel.

OpenStudy (anonymous):

So, looking at the first piece specifically, what is going to cancel? \[\frac{(x-4)(x-2)}{1}\frac{x}{x-2}\]

OpenStudy (anonymous):

I cancelled in my head xD thats my favorite part of this once i ACTUALLY get to this part without making a mistake :P

OpenStudy (anonymous):

Hey - even better! :)

OpenStudy (anonymous):

so you would cance the (x-2)

OpenStudy (anonymous):

Exactly!

OpenStudy (anonymous):

then your left with (x-4)x/1 right?

OpenStudy (anonymous):

Yes, although because the denominator is 1, we don't need to write it anymore.

OpenStudy (anonymous):

So then because you were able to cancel in your head, you see how we're left with: \[x(x-4)+1(x-2)=2\]

OpenStudy (anonymous):

The canceling is the payoff for picking the right LCD. If you've done your job right, at this point, there should be no fractions left.

OpenStudy (anonymous):

oh yah forgot that second piece :P

OpenStudy (anonymous):

Oh, be sure to remember all of them. We have 3 pieces to this equation. :)

OpenStudy (anonymous):

So from here, what do we do?

OpenStudy (anonymous):

Bring the equation together?

OpenStudy (anonymous):

That's right - distribute, combine like terms... What will that leave you with?

OpenStudy (anonymous):

(x-4)((x-2)(x-2)

OpenStudy (anonymous):

I'm not sure what you did there.

OpenStudy (anonymous):

Here's where you want to go with that:

OpenStudy (anonymous):

First, use the distributive properto to get rid of those parenthesis:\[x(x-4)+1(x-2)=2\] Then combine the -4x with the +x\[x^2-4x+x-2=2\] Now, recognize that this thing is going to end up quadratic, so we want to set it equal to zero. \[x^2-3x-2=2\]You can subtract 2 from both sides to do that. \[x^2-3x-4=0\]

OpenStudy (anonymous):

WOW!

OpenStudy (anonymous):

Do you need me to explain any of those steps a little better?

OpenStudy (anonymous):

I think that was the best you could explain mate xD

OpenStudy (anonymous):

Excellent!

OpenStudy (anonymous):

Alright, so we're left looking at this: \[x^2-3x-4=0\] Now you want to factor that. What does it leave you with?

OpenStudy (anonymous):

(x-4)(x+3) x=4 x= -3 ?

OpenStudy (anonymous):

Well, it can't be (x-4)(x-3) because if you tried to FOIL that out, you'd get: \[x^2-7x+12\]

OpenStudy (anonymous):

We're looking for numbers that multiply to -4 and add to -3

OpenStudy (anonymous):

Oops I meant (x-4)(x+1)

OpenStudy (anonymous):

That's right! That's perfect! :)

OpenStudy (anonymous):

So we're left with:\[(x-4)(x+1)=0\]

OpenStudy (anonymous):

Now we can use the Zero-Product Property: \[x-4=0\]&\[x+1=0\]

OpenStudy (anonymous):

This is GREAT xD

OpenStudy (anonymous):

So, what do you get for the answers?

OpenStudy (anonymous):

x = 4 and x = -1 ?

OpenStudy (anonymous):

There's a little trick to this one. Because these are rational equations, we have to check to see if those are good answers. According to the equation, those are our answers, but they may turn out to be bad x's. We have to plug them in to the original equation and make sure they don't cause division by zero.

OpenStudy (anonymous):

Only one of those answers is correct. The other is a bad x.

OpenStudy (anonymous):

So, my question for you, which is the right one, and which is the bad one?

OpenStudy (anonymous):

how would i know D: lol

OpenStudy (anonymous):

Well, only one of them will cause division by zero with our original equation.

OpenStudy (anonymous):

am i suppsed to plug it in and find out?

OpenStudy (anonymous):

You could plug them in and find out that way, or you can just eyeball it. Have a look: When we factored the original equation we were left with this:\[\frac{x}{x-2}+\frac{1}{x-4}=\frac{2}{(x-4)(x-2)}\]

OpenStudy (anonymous):

Our two solutions are: x=4 and x=-1

OpenStudy (anonymous):

Now, the -1 won't cause any problems. But that 4 will because look at the second fraction. If we plug a 4 in for x, what does the bottom become?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Exactly - and we can't ever divide by 0. So that makes x=4 a "bad x". Because x=4 causes division by 0 in our original equation, we have to throw it out. So this question only has 1 answer. x=-1 :)

OpenStudy (anonymous):

So you always want to check your answers for these to make sure they're not bad x's.

OpenStudy (anonymous):

Sometimes there are no bad x's, sometimes (like this time) there is 1 bad x, and other times, both solutions are bad x's.

OpenStudy (anonymous):

And that's how they're done! :) I hope that helped!!

OpenStudy (anonymous):

THANKS!! xD

OpenStudy (anonymous):

You're welcome!!